# Prioritizing Security Practice Adoption: Empirical Insights on Software Security Outcomes in the npm Ecosystem

### v1

@@ 5.1.1. Regression Model:
We selected Generalized Linear Models (GLMs) (nelder1972generalized, ), given the non-normal distribution of our dataset and our target variables are count data. GLMs are specifically designed for count data and provide interpretable coefficients. Coefficients quantify the expected change in the target variable for a unit increase in a predictor. To identify the best-fitting GLM, we systematically evaluated the dataset characteristics using a data-driven approach. Our selection process was guided by two factors: the Overdispersion Test (variance greater than the mean) and the zero-inflation test(a large proportion of zero values in target variables). We first started with the Poisson Regression (hausman1984econometric, ) as our baseline model and found that the dataset contains overdispersion for Vul_Count , MTTR, and MTTU, and zero inflation exists for Vul_Count only. We selected Negative Binomial (NB) Regression (hilbe2011negative, ) for MTTR and MTTU to address overdispersion and Zero-Inflated Negative Binomial (ZINB) (ridout2001score, ) to handle both overdispersion and zero inflation for Vul_Count. Since GLMs inherently apply a log link function instead of log transformation, we used the raw target variables to maintain the interpretability of the coefficients.
5.1.2. Causal Analysis:
While regression models can identify associations between “aggregated Scorecard score” and security outcomes, they do not establish causality due to the potential influence of confounding factors. For example, the predictors aggregated Scorecard score and control are correlated, which could bias the estimation of the true impact of security practices on security outcomes. To address this, we used Propensity Score Matching (PSM) (austin2009balance, ), a statistical technique that simulates a randomized experiment by pairing repositories with similar characteristics. PSM matched the treated units (repositories with aggregated Scorecard scores above the median) with control units (below the median) based on covariates. Covariates are control variables that could influence both aggregated Scorecard scores and security outcomes. Using logistic regression, we estimated each repository’s likelihood of having a high aggregated Scorecard score. We then applied nearest-neighbor matching with a caliper adjustment(0.1) to pair treated repositories with control repositories based on similar covariates. The similarity of matching was determined using propensity scores, which represent the likelihood that a repository belongs to the treated group based on the repository characteristics. To validate the matching process, we used standardized mean differences (SMD) to assess balance across covariates. Finally, we calculated the Average Treatment Effect (ATE) using an inverse probability-weighted (IPW) estimate (ATE weighted) that adjusts for any residual differences. Computing the Average Treatment Effect (ATE), we estimated how much a package’s security outcome would change if it had a higher aggregated Scorecard score.
@@ 5.2. RQ1: Result
Figure 3 shows a sharp decline in security outcomes (in Vul_Count, MTTR, and MTTU) as the aggregated Scorecard score increases. The highest bin (7.3–9.1) reflects the top-scoring repositories, as no project achieved a perfect score. However, the plot only reflects an overall trend and does not account for confounding factors and associations. To formally assess the relationship, we use regression analysis and propensity score matching (PSM). Table 3 demonstrates that a higher aggregated Scorecard score is associated with improved security outcomes (negative coefficient). In the ZINB model, a coefficient of –1.4 for Vul_Count implies that a unit increase in the aggregated Scorecard score corresponds to a 1.4 count decrease in the expected number of vulnerabilities. The finding is also supported by negative ATE estimates (–1.9) in the PSM analysis. For MTTU, both the NB model (coefficient –0.2) and PSM estimates (–0.5) indicate that repositories with higher aggregated Scorecard scores update dependencies more rapidly. By contrast, the NB model for MTTR shows a downward trend (Figure 3) and a negative coefficient (–0.1). However, the PSM analysis yields a positive overall ATE (74.7), indicating that, when controlling for control variables, repositories with higher aggregated Scorecard scores tend to have longer update delays for vulnerable components.
Table 3. GLMs and PSM analysis between Aggregated Scorecard Score and Security Outcomes
Dep. Var Model Coef. p-value ATE(Weighted)
Vul_Count ZINB -1.4 ¡0.001 -1.9
MTTR NB -0.1 ¡0.001 74.7
MTTU NB -0.2 ¡0.001 -0.5
We conducted a segment-based analysis for MTTR to investigate further the discrepancy by stratifying the dataset into subgroups based on repository characteristics using a median split. The segment analysis revealed that projects in the upper groups for size, contributors, dependency count, and downloads and the lower group for age tend to exhibit negative ATEs (indicating faster MTTR). In contrast, repositories in the lower groups for size, contributors, dependency count, and downloads and in the upper age group show positive ATEs (indicating slower MTTR). Overall, the segment analysis reveals that the effectiveness of security practices adoption in reducing MTTR is context-dependent. Larger, newer repositories with more contributors, dependencies, and downloads benefit from adopting security practices, resulting in faster MTTR. In contrast, smaller but older repositories with fewer contributors, downloads, and dependencies may experience increased overhead from adopting security practices, leading to longer MTTR.
Summary: Our analysis shows that higher aggregated Scorecard scores are associated with fewer vulnerabilities and shorter MTTU. However, while regression models indicate shorter MTTR, causal analysis reveals that repositories with higher aggregated Scorecard scores may take longer to update vulnerable components since project characteristics also influence the relationship.
Table 4. Distribution of Security Practices Adoption
Metric % -1 % 0 % 1-10
Branch Protection 0.005% 85.9% 14%
Code Review 1.4% 74.6% 24%
License 0 31% 69%
Maintained 0.001% 86.7% 13.3%
Pinned Dependencies 72.2% 0 27.9%
@@
Table 5. R2, Security practice Importance Ranked and PSM Insights
Dep. Var R2 for RF Ranked Security Practices PSM Effect (ATE weighted direction)
@@
Vul_Count 0.38 Code Review, Maintained, Branch Protection, License, Pinned-Dependencies Code Review (+), Others (-)
@@
MTTR 0.42 Maintained, Pinned-Dependencies, Branch Protection, Code Review, License All (+)
@@
MTTU 0.2 Maintained, Code Review, Pinned-Dependencies, Branch Protection, Security Policy Code Review (+), Branch Protection(+), Others (-)
6. RQ2: Associations between security practices and outcomes
@@
In this section, we discussed our RQ2: Which security practices are most strongly associated with security outcomes? Section 6.1 discusses our methodology to answer our RQ2, and Section 6.2 discusses our result.
@@ 6.1. RQ2: Methodology
For RQ2, we investigate the association between individual security practices and security outcomes. Unlike RQ1, where we analyzed aggregated Scorecard scores as a predictor, for RQ2, we analyzed twelve distinct security practices as predictors. The large number of predictors, non-normal distribution, and non-linear relationships among variables (section 4) introduce additional complexity to the dataset. We selected a tree-based regression model as these models are well-fitted for datasets with complex, non-linear relationships and high-dimensional feature interactions. We also conducted the feature importance ranked analysis to identify the security practices that most strongly contribute to predicting security outcomes. We performed causal analysis to explore potential causal relationships since regression and feature importance identify associations but do not establish causality.
@@ 6.1.1. Regression Model:
Among tree-based models, we chose Random Forest (RF) (dewi2019random, ) due to its ensemble learning approach, which aggregates multiple decision trees to improve model performance and reduce overfitting. Unlike single decision trees, which can be highly sensitive to variations in training data, RF mitigates the issue by averaging predictions across multiple trees. We also utilized Gini Importance (Mean Decrease in Impurity) as a feature importance technique to quantify each security practice’s contribution toward reducing prediction error. Gini Importance measures a predictor’s contribution by assessing how often the predictor is used for splitting the data in the decision trees of the RF model, highlighting the practice’s contribution in accurately predicting security outcomes. We first preprocessed our data, converting categorical security practices into appropriate numerical or binary formats (section 4.1). We then split the dataset into training (70%) and testing (30%) subsets. The model was trained on the training dataset using default hyperparameters. Five-fold cross-validation was applied to estimate and reduce bias, calculating mean squared error (MSE) and R² scores. The final model’s performance and feature importance scores were validated on the reserved testing set (Table 5).
6.1.2. Causal Analysis
We followed the methodology outlined in Section 5.1.2 for our causal analysis. We conducted the PSM analysis only on the top 5 security practices ranked by the Gini feature importance measure. The top 5 was selected because, after that, the feature importance value was almost near 0, indicating those practices did not have a strong association with security outcomes. The aggregated Scorecard score is computed as a weighted average of 18 individual security practices; only 25 repositories resulted in a score of zero. On the contrary, individual security practices exhibit a higher rate of non-adoption or security misconfiguration (score -1, score 0), as shown in Table 4 (top 5 based on feature ranked). The disparity occurs because aggregation combines multiple practices, ensuring a nonzero score even when some practices are not adopted. In contrast, an individual practice remains at 0 or -1 if the practice is not implemented. As a result, individual security practices had a higher proportion of repositories with a score of 0 or -1, leading to a highly imbalanced dataset for PSM analysis. To mitigate imbalance and ensure balanced treatment and control groups, we applied a sampling approach for the top 5 security practices. We first included all repositories with scores between 1 and 10, then randomly sampled an equal number of repositories with a practice score of zero to create a representative baseline. We then created a binary treatment variable by splitting the combined dataset based on an adaptive threshold (below and above the median score). However, for Maintained and License metrics, the median score was equal to the maximum observed score (indicating minimal variance), and we adjusted the threshold to the midpoint between the minimum and maximum scores to ensure separation between high and low adoption groups. After constructing balanced groups, we followed the technique discussed in Section 5.1.2, where we used nearest-neighbor matching with a caliper adjustment and computed the ATE score (ATE weighted).
@@ 6.2. RQ2: Result
Our analysis employed RF regression, complemented by feature importance analysis and Propensity Score Matching (PSM), to identify security practices most strongly associated with security outcomes. Table  5 summarizes our findings, illustrating the variance explained (R²) by security practices and each practice’s directional effects based on the PSM analysis’s ATE-weighted values. The R² value indicates that security practices collectively explain a moderate portion of the variance in Vul_Count (R² = 0.38) and MTTR (R² = 0.41), and a relatively smaller proportion for MTTU (R² = 0.2). Notably, the lower R² for MTTU implies that additional external factors or project-specific characteristics beyond those captured in our current model likely contribute to MTTU. From the feature importance analysis, Code Review, Maintained, Pinned Dependencies, and Branch Protection practices consistently rank highly across all three outcome metrics. License considerations showed relevance for both Vul_Count and MTTR, whereas Security Policy appears influential for MTTU, highlighting the variability of practice impacts across distinct outcome measures.
In the PSM analysis of individual security practices, we did not consistently observe a negative (i.e., improving) effect across all security outcomes. Most security practices show a negative ATE for Vul_Count and MTTU (Table  5), implying fewer vulnerabilities and faster update frequency. However, Code Review consistently exhibits a positive relationship with all three security outcomes, suggesting that the practice aids in detecting more vulnerabilities. As emphasized in CISA’s Secure by Design guidance (cisa2023securebydesign, ), a rise in reported vulnerabilities can reflect a healthy security posture due to code review and testing by the community. For MTTR, all practices exhibit a positive ATE, indicating longer remediation times for repositories adopting these practices. Figure 3 shows that 90% npm repositories fall into low-security score bins (0.0–3.6), where MTTR is high (1200 days), suggesting these projects do not prioritize dependency security or vulnerability remediation. Table 4 further highlights that many npm packages lack security practices adoption, reinforcing this trend. While higher aggregated security scores and individual practices are associated with fewer Vul_Count and shorter MTTU, MTTR remains positive for both cases, suggesting MTTR is influenced by factors beyond security practice adoption, such as project maintainability, resource availability, and dependency complexity. Unlike RQ1, where segmentation revealed how repository characteristics impact MTTR, we could not perform a similar analysis in RQ2 due to data imbalance (Section 6.1). The limited sample size in the PSM analysis further constrained our ability to isolate project-specific influences on MTTR. Therefore, we cannot conclusively confirm whether the observed increase in MTTR consistently indicates slower remediation or whether the effect might vary depending on project-specific characteristics, maintainability, or dependency complexity. Overall, Vul_Count , MTTR, and MTTU exhibit similar trends as security outcomes metrics. However, MTTR is more complex and may have an influence on external factors such as project maintainability, resources, and dependency complexity that need further research.
Summary: Our analysis shows that Code Review, Maintained, Pinned Dependencies, and Branch Protection are strongly associated with security outcomes. While individual security practice is associated with fewer vulnerabilities and faster dependency updates, MTTR remains high, suggesting that factors beyond security practice adoption, such as project maintainability, resources, and dependency complexity, influence MTTR.

### v2

@@ 5.1.2. Regression Model:
At this stage, we focus on interpreting how security outcomes change per unit change in the aggregated Scorecard score (e.g., a coefficient of -0.1 implies a 10% decrease in the outcome for every one-unit increase in aggregated Scorecard score). We selected Generalized Linear Models (GLMs) (nelder1972generalized, ), given the non-normal distribution of our dataset, and our target variables are count data. GLMs are explicitly designed for count data and provide interpretable coefficients. Coefficients quantify the expected change in the target variable for a unit increase in a predictor. To identify the best-fitting GLM, we systematically evaluated dataset characteristics using standard diagnostic tests: overdispersion (variance $<$ mean) and zero-inflation. We began with Poisson Regression (hausman1984econometric, ) as our baseline model and found significant overdispersion across all outcomes, with zero-inflation present only in Vul_Count. We applied Negative Binomial (NB) Regression (hilbe2011negative, ) to address overdispersion and attempted Zero-Inflated Negative Binomial (ZINB) (ridout2001score, ) for Vul_Count to handle both issues simultaneously. However, ZINB models encountered numerical convergence failures with precision loss warnings during maximum likelihood estimation. Therefore, we applied NB regression across all outcomes, providing stable parameter estimation while adequately addressing overdispersion.
5.1.3. High vs. Low Aggregated Scorecard Scores: Comparing Similar Repositories
To move beyond correlational evidence and establish causality, we employed PSM to simulate experimental conditions by comparing high and low-scoring repositories with similar project characteristics. PSM (austin2009balance, ) is a statistical technique that simulates the conditions of a randomized experiment by matching repositories with similar characteristics but differing in their level of security practice adoption. Similar project characteristics refer to repositories matched on the same covariates used as control variables in section 3.1.3.
Repositories are classified as “treated” if their aggregated scores fall within the top quartile, with the remaining projects serving as a control group. We tested multiple treatment thresholds (median, tertile, quartile) and selected the top quartile ($\geq$ 3 on a 1-10 scale) to ensure sufficient contrast between treatment and control groups while maintaining adequate sample sizes. Using logistic regression on covariates that influence both aggregated scores and security outcomes, we estimated each repository’s propensity score (probability of having a high aggregated score given its characteristics), then matched treated and control repositories with similar propensities. We then applied nearest-neighbor matching with a 0.1 standard deviation caliper to pair treated and control repositories with similar propensities. We validated matching quality using standardized mean differences (SMD) and calculated Average Treatment Effects (ATE) through difference-in-means with bootstrap confidence intervals (CI) to quantify uncertainty. We complement this analysis with NB on post-matched samples (containing only the matched treated and control repositories).
Table 2. Segment-specific ATEs and percentage impacts. Strongest tercile per variable shown.
Outcome Controls Segment ATE Impact
Vul_Count Contributors_CT Many -8.2 28.9%
Age Mature -11.0 42.2%
Commit Staleness Fresh -9.4 47.5%
Downloads High -9.3 35.9%
Size Large -8.4 27.7%
Dependency Medium-Dep -9.8 37.0%
MTTR Contributors_CT Many -350.1 41.7%
Age Mature -429.8 37.5%
Commit Staleness Fresh -244.8 37.3%
Downloads High -269.5 40.5%
Size Large -315.0 40.2%
Dependency Few -245.7 28.1%
MTTU Contributors_CT Many -87.2 19.1%
Age Mature -86.0 12.8%
Size Small 146.3 40.0%
Commit Staleness Stale 212.1 23.5%
Downloads Few 146.1 38.5%
Dependency High-Dep -81.0 20.6%
5.1.4. Impact Across Segments of Each Repository Characteristic
While our PSM analysis establishes an overall causal effect by controlling for repository characteristics, it estimates one average treatment effect without examining how this effect varies across different repository types. For instance, the security benefits of higher aggregated scores might be stronger for large repositories compared to small ones. To examine whether this effect differs in magnitude across different levels of repository characteristics, we conducted post-hoc heterogeneity analyses on the matched sample. We stratified the matched repositories into terciles based on each control variable using quantile-based binning and calculated segment-specific ATEs, comparing high and low scorecard scores within each segment. This approach identifies which repository contexts benefit most from higher aggregated scores. We used bootstrap CI to quantify uncertainty around segment-specific ATEs.
@@ 5.2. RQ1: Result
Table 3. Combined GLM and PSM analysis for RQ1
Outcome NB Coef NB Coef (PSM) PSM ATE (95% CI)
Vul_Count -0.20 -0.26 -5.2 [-5.7, -4.7]
MTTR -0.05 -0.16 -216.8 [-241.8,-193.2]
MTTU 0.03 -0.04 -52.3 [-60.2, -44.8]
Data Exploration: Figure 2 shows a sharp decline in security outcomes (indicating improvement) as aggregated Scorecard scores increase. The highest bin (7.3–9.1) reflects the top-scoring repositories, as no project achieved a perfect score. However, the plot only reflects an overall trend and does not account for confounding factors and associations.
Quantitative Confirmation: Table 3 confirms visual trends (Figure 2) through three analytical approaches: full-sample regression, PSM post-matched sample regression, and PSM causal analysis. The columns show full-sample NB coefficients, post-matching NB coefficients, and post-matching ATEs with 95% confidence intervals. For Vul_Count, both full-sample (–0.20) and post-matching (–0.26) NB models show negative coefficients, with ATE indicating 5.2 fewer vulnerabilities. MTTR demonstrates consistent improvements: full-sample (–0.05), post-matching (–0.16), and ATE (–216.8 days). For MTTU, while the full-sample model shows a positive coefficient (0.03), post-matching (–0.04) and ATE (–52.3 days) reveal faster updates, suggesting confounding factors mask the relationship.
Repository Context-Dependent Effects: While our PSM analysis established overall causal effects, we conducted segment-specific analyses to understand where these effects are strongest. Table 2 presents the most pronounced effects within each repository characteristic, with the ”Segment” column identifying the specific subgroup and the ”Impact” column quantifying percentage improvements relative to the control baseline within that segment.
Across all segments, Vul_Count consistently shows positive security outcomes (i.e., negative ATE values), with improvements ranging from 27.7% to 47.5%. The most substantial reductions in open vulnerabilities are observed in mature repositories with fresh commit staleness (47.5% and 42.2%, respectively), indicating that actively maintained and mature repositories with high aggregated scores tend to have fewer vulnerabilities. Medium-dependency group repositories show the greatest vulnerability reduction, highlighting the effect of dependency complexity. MTTR improvements are also observed across all segments. Mature repositories have the largest absolute improvement (429.8 days faster), followed by those with many contributors (350.1 days). Repositories with fewer dependencies and fresh commit staleness show meaningful reductions as well (-245.7 and -244.8 days). Then, repositories with higher download counts tend to have fewer Vul_Count and faster MTTR. MTTU reveals the most heterogeneous pattern, with some segments showing improvements while others exhibit deterioration. Small repositories by size and download and repositories with stale commit activity show increases in MTTU, suggesting that top quartile aggregated score repositories that are small, less downloaded, or less active have slower dependency updates compared to low Scorecard score repositories in the same segment category. Conversely, repositories with many contributors, mature age, and high dependency counts demonstrate faster updates.
Summary of RQ1 Findings:
  • Overall Effects: Higher aggregated Scorecard scores improve security outcomes, with repositories achieving 5.2 fewer vulnerabilities, 216.8 days faster MTTR, and 52.3 days faster MTTU.
  • Repository Context Matters: The adoption of aggregated security practices improves security outcomes. The segment-specific analysis reveals nuanced patterns in how repository characteristics influence security practice effectiveness.
  • Organizational Capacity: Mature repositories with many contributors consistently show better outcomes, suggesting that well-staffed and mature projects with high adoption of security practices benefit from improved security outcomes.
  • Project size, activity, and downloads: Project size, recent commit activity, and download volume are associated with improvements in Vul_Count and MTTR. However, their absence creates challenges (i.e., smaller, inactive, or less-downloaded projects) and correlates with slower MTTU, highlighting how inactivity and smaller projects with fewer downloads can hinder general updates.
  • Dependency Complexity: Different dependency levels excel in different areas. Medium-dependency projects have the fewest open vulnerabilities, likely balancing manageable complexity. Few-dependency projects remediate issues fastest, benefiting from simpler, more transparent dependency trees. Large-dependency projects excel at general updates, likely leveraging automation to manage scale effectively.
While selecting dependencies, organizations should prioritize mature, actively maintained large projects with many contributors, fewer dependencies, and high downloaded packages, as our analysis showed that such projects have fewer vulnerabilities and faster MTTR due to better security performance. Conversely, smaller, inactive packages with fewer downloads require enhanced monitoring and proactive security assessments to compensate for slower MTTU.
@@
Table 4. R2, Security practice Importance Ranked and PSM Insights. Positive ATE indicates increased outcome (worse), negative ATE indicates reduced outcome (better).
Outcome R2 for RF Ranked Security Practices PSM Effect (ATE direction)
@@
Vul_Count 0.39 Contributors, Code Review, License, Pinned-Dependencies, Branch Protection, CI-Tests, Dependency-Update-Tool Code Review (+), CI-Tests (mixed), Branch Protection (+), Dependency-Update-Tool (+), Others (-)
@@
MTTR 0.50 Contributors, Code Review, License, CI-Tests, Pinned-Dependencies Code Review (+), Others (-)
@@
MTTU 0.32 Contributors, License, Code Review, CI-Tests, Pinned-Dependencies Code Review(+), CI-Tests (mixed), Others (-)
6. RQ2: Impact of individual security practices and outcomes
@@
In this section, we discussed our RQ2: Which security practices are most strongly associated with security outcomes? Section 6.1 discusses the methodology for RQ2, and Section 6.2 discusses the result.
@@ 6.1. RQ2: Methodology
For RQ2, we investigate the association between individual security practices and security outcomes. Unlike RQ1, where we analyzed aggregated Scorecard scores as a predictor, for RQ2, we analyzed eleven security practices as predictors after exclusion (Section 3.2.3 and 4.2) alongside six control variables. Due to the large number of predictors, non-normal distribution, and non-linear relationships (Section 4), we selected a tree-based regression model suited for handling complex, high-dimensional data. We also conducted the feature importance ranked analysis to identify the security practices that most strongly contribute to predicting security outcomes. We performed causal analysis to explore potential causal relationships since regression and feature importance identify associations but do not establish causality.
@@ 6.1.1. Regression Model:
We employed Random Forest (RF) regression due to its robustness to non-linear relationships and ability to handle high-dimensional data without overfitting. Following preprocessing (section 4.1), the dataset was split into training (70%) and testing (30%) subsets using random sampling. The model was trained on the training dataset using hyperparameters (n_estimators=100, random_state=42). To ensure robust performance estimation and reduce bias, five-fold cross-validation was applied exclusively to the training data, calculating R² scores(Table 4).
6.1.2. Security Practice Ranking
: Feature importance was quantified using Gini Importance, extracted directly from the trained Random Forest model’s feature importances attribute, which measures each feature’s contribution to reducing node impurity across all decision trees in the ensemble. Features were ranked by their grouped importance scores using dense ranking methodology (method= dense) within each dependent variable. Since the preprocessing step decomposed high-missing features (CI-Tests, Pinned-Dependencies) into separate components, we grouped all feature importance scores by their base security practice to provide more interpretable ranking results. The method involved summing the decomposed components for CI-Tests and Pinned-Dependencies, while non-decomposed features retained their individual importance scores. Using the ranked importance scores of 11 practice metrics, we selected the top-performing security practices that collectively accounted for approximately 80% of the cumulative feature importance across all three outcome metrics (ucla2023efa, ). The approach provides interpretable insights into which security practices have the strongest overall predictive power for each outcome variable (Table 4).
6.1.3. High vs. Low Security Practice Scores: Comparing Similar Repositories
We conducted PSM analysis on the top-ranked security practices from Table  4. We focused on these top practices because the remaining ones demonstrated negligible importance values while the selected top features collectively accounted for approximately 80% of the cumulative feature importance. However, individual security practices exhibit extreme binary distributions, with most repositories scoring either 0 (not adopted) or 10 (fully implemented), creating an extreme class imbalance that violates PSM overlap assumptions. Unlike aggregated Scorecard scores, which combine multiple practices to ensure nonzero values, individual practices remain at their adoption extremes. To enable valid causal inference, we applied balanced sampling to create sufficient propensity score overlap. Our PSM methodology employed a universal threshold of 5.0, classifying repositories with scores above this threshold as treated units and those at or below as control units. Following balanced sampling, after constructing balanced groups, we followed the technique discussed in Section 5.1.3, where we used nearest-neighbor matching with a caliper adjustment and computed ATE.
Unlike RQ1, where segmentation revealed how repository characteristics impact outcome, we could not perform a similar analysis in RQ2 due to data imbalance (Section 6.1). The limited sample size in the PSM analysis further constrained our ability to isolate project-specific influences on security outcome metrics.
@@ 6.2. RQ2: Result
Quantitative Confirmation: Table  4 summarizes our RQ2 findings, analyzing individual security practices’ relationships with security outcomes. Security practices collectively explain a moderate portion of variance across outcomes: R² = 0.39 for Vul_Count , R² = 0.50 for MTTR, and R² = 0.32 for MTTU. The higher explanatory power for MTTR suggests stronger relationships between security practices and MTTR compared to Vul_Count or MTTU.
Important Security Practices: The analysis reveals consistent patterns across outcome metrics. Contributors emerge as the most influential factor across all three outcome metrics, followed by Code Review and License practices. CI-Tests and Pinned-Dependencies consistently rank among the top practices. The ranking pattern suggests that collaborative human factors (Contributors) and development workflow practices (Code Review, CI-Tests) have an influential, measurable impact on security outcomes. Branch Protection and Dependency Update Tool ranks among the top seven features required to reach the 80% cumulative importance threshold for Vul_Count. In contrast, MTTR and MTTU reach the threshold with five features, but including these two additional practices raises the cumulative importance to nearly 92% for MTTR and MTTU, indicating these practices also have a stronger impact on the outcomes.
Impact of Security Practice: The PSM analysis reveals complex and counterintuitive relationships between individual security practices and outcomes. Code Review consistently exhibits a positive ATE across all three security outcomes, indicating associations with more vulnerabilities, slower MTTR, and MTTU. This paradoxical finding aligns with CISA’s Secure by Design guidance (cisa2023securebydesign, ), a rise in reported vulnerabilities can reflect a healthy security posture due to code review and testing by the community. CI-Tests show a mixed pattern: projects with pull-request (PR) infrastructure tend to have faster MTTR but also higher vulnerability counts and slower MTTU, possibly due to added complexity from collaboration. However, projects with PR infrastructure that also achieve high CI-Test scores consistently perform better across all security metrics. The finding suggests that PR workflows alone are insufficient; security benefits arise when effective CI practices are fully implemented. Beyond Code Review and CI-Tests, other practices show more intuitive associations. Contributors, License, and Pinned Dependencies are associated with negative ATEs, suggesting improved outcomes if adopted by repositories. For Vul_Count, the Branch Protection and Dependency Update Tool shows positive ATEs, suggesting these practices enhance vulnerability detection capabilities.
Our analysis shows that Contributors, Code Review, License, CI-Tests, and Pinned-Dependencies are the most influential security practices across all outcome metrics. While adoption of Code Review practice shows positive associations with more vulnerabilities and slower MTTR and MTTU, Contributors, License, and Pinned Dependencies demonstrate negative ATEs, indicating improved security outcomes. CI-Tests show mixed effects depending on CI-test adoption.